OpenStudy (anonymous):

Can someone please help & explain this question? How many solutions does the system have? x = -4y + 4 2x + 8y = 8

OpenStudy (anonymous):

x = -4y + 4 --- (1) 2x + 8y = 8 --- (2) there are a few ways to solve this :3 you can: 1. substitution plug in (1) to (2) 2 (-4y + 4) + 8y = 8 try working it out from there :p 2. elimination get (3) from (1) x + 4y = 4 --- (3) then times it by 2 so you can eliminate it by (2) 2x + 8y = 8 --- (3) 2x + 8y = 8 --- (2) ------------------- - ..............wait. something is wrong with this question > 3> 2 (-4y + 4) + 8y = 8 -8y + 8 + 8y = 8 8 = 8 ....................

OpenStudy (anonymous):

idk get someone pro to answer this sorry xDD

OpenStudy (across):

x = -4y + 4 2x + 8y = 8 x = -4y + 4 2(x) = 2(-4y + 4) one equation is a multiple of the other. that means you have infinitely many solutions

OpenStudy (anonymous):

thanks you so much!